OpenStudy (anonymous):

(4, -5) and that passes through (-1, 3)?

OpenStudy (anonymous):

I got 1/6 = a, but I don't know how it's final answer is f(x) = 1/6(x^ - 7)

Directrix (directrix):

What are you finding with the two points? Slope? Equation of a line?

OpenStudy (anonymous):

equation of the parabola

OpenStudy (anonymous):

with x- intercepts

Directrix (directrix):

Is it possible for just two points to determine a parabola? Are these special points, maybe the vertex and some other point on the parabola. I don't recognize this < f(x) = 1/6(x^ - 7) >> as the equation of a parabola.

Directrix (directrix):

Do you have instructions for the problem that you could post here?

OpenStudy (anonymous):

well this is what I did f(x) = a(x-r)(x-s) f(-3) = a(1-4)(1+5) -3 = a(-3)(6) -3/18 = -18/18 a 1/6 = a

OpenStudy (lalaly):

can u post the whole question?

OpenStudy (anonymous):

Determine the equation of the parabola with x- intercepts.

OpenStudy (lalaly):

x intercepts 4 and -5?